OpenStudy (anonymous):

Find the center of the circle defined by the equation x^2+y^2-4x-4y+4=0

OpenStudy (anonymous):

You want to complete the square for x and y. \[x^2-4x+(4-4)+y^2-4y+(4-4)+4=(x-2)^2+(y-2)^2=4\]

OpenStudy (anonymous):

So your center is (2,2)

OpenStudy (anonymous):

^ was going to write the exact thing lol

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Thanks love!

OpenStudy (anonymous):

No problem :P